OpenStudy (anonymous):

Line Integral help please! r=(5costsint)i+(5sin^2t)j+(5cost)k this curve lies on a sphere with equation x^2+y^2+z^2=25 Evaluate the c integral z ds using this curve. I will write more exact below.

OpenStudy (anonymous):

Evaluate \[\int\limits_{c}^{} z ds\] for curve above.

OpenStudy (anonymous):

t from where to where.

OpenStudy (anonymous):

I think 0 to 2pi because its a sphere?

OpenStudy (anonymous):

well first x^2+y^2+z^2=25 Solve for z take z and replace x by 5costsint, y by 5sin^2sin and k by 5cost Basically plug into x,y, and z the i,j ant components this gives z(t) (z with respect to t) then take tha magninuted of vector | r | I assume you know how to do that then take that magnitude and multiply by z(t): |r|*z(t) Finally integrate wrt: intigrate (|r|*z(t)*dt) You assume from 0 to 2pi

OpenStudy (anonymous):

\[ |r| =5\\ \int_0^{2 \pi} 5 ( 5 \cos(t)) dt=0 \]